OpenStudy (anonymous):

Help Please Of 450 college students, 120 are enrolled in math, 225 are enrolled in English, and 55 are enrolled in both. If a student is selected at random, find the probability of the following. (Round your answers to three decimal places.) The student is not enrolled in English or is enrolled in mathematics.

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

draw a venn diagram

OpenStudy (anonymous):

I added 120+225-55=290 and divided 290/450 but i still keep getting the wrong answer am i doing something wrong?

OpenStudy (anonymous):

299 is are those students who either take engish or math. You're supposed to find those who do not take English or do not take math. P(not E or not M)

OpenStudy (anonymous):

typo. *290 are those...*

OpenStudy (anonymous):

If its neither would i subtract?

OpenStudy (anonymous):

well, you're doing P(not English or not English) yes?

OpenStudy (anonymous):

Yes Not in English or math

OpenStudy (anonymous):

what is P(A or B) = ?

OpenStudy (anonymous):

Is it 120+225=345?

OpenStudy (anonymous):

nope. Just finish that formula for me

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm so lost 0.0

OpenStudy (anonymous):

P(A or B) = P(A) - P(B) - P(A and B). Look familiar?

OpenStudy (anonymous):

120-225-55?

OpenStudy (anonymous):

nope. what is P(not English or not Math) = ? using the formula above?

OpenStudy (anonymous):

Alrighty,Thanks =)